After 4 years of investment you earned $252.00 in simple interest. The annual rate of interest was 5.25%. How much wes your initial investment (principal)?
Answer by Theo(13342)

simple intrest assumes no compounding.

the formula is i = p * r * n

i = interest
p is the principal
r is the interest rate per time period.
n is the number of time periods.

your annual intrest rate is 5.25% = .0525.
your number of years is 4.
your interest is 252.

the formula becomes 252 = p * .0525 * 4

divide both sidesof this equation by (.0525 * 4) and you get:

252 / (.0525 * 3) = p

solve for p to get p = 1200.

you invested 1200 at .0525 interest rate per year for 4 years.

.0525 * 1200 = 63 * 4 = 252.

solution looks good.

original investment is 1200.